A certain lottery has 2828 numbers. In how many different ways can 44 of the numbers be​ selected?
Answer by stanbon(75887)

A certain lottery has 28 numbers. In how many different ways can 4 of the numbers be​ selected?
28C4 = (28*27*26*25)/(1*2*3*4) = 7*9*13*25 = 20,475 ways
-----------
Cheers,
Stan H.
---------